A 5.5 kg rifle fires an 11g bullet with a velocity of 1000 m/s.

(a)   What is the recoil velocity of the rifle?

(b)  If the shooter holds the rifle firmly against his shoulder, what will be the recoil velocity? Assume the mass of the shooter is 120 kg.


1
Expert's answer
2021-03-21T11:33:52-0400

a)


"Mv=mV\\5.5v=(0.011)(1000)\\v=2\frac{m}{s}"

b)


"M'v'=mV\\(120+5.5)v'=(0.011)(1000)\\v'=0.088\frac{m}{s}"